OpenStudy (anonymous):

A kilogram is about 2.2 pounds. A cube of water 30 cm on edge weighs about how many pounds?

OpenStudy (anonymous):

We know the density of water to be \(1 g/cm^3\)From dimensional analysis, we can observe that\[\rho \left [g \over cm^3 \right] \cdot V \left [ cm^3 \right] = m [g]\] Therefore, we need to calculate the volume of the cube, which is simply \(V = a^3\), and multiply it by the density.

OpenStudy (anonymous):

the volume is 27L i figured that out...but now what?

OpenStudy (anonymous):

Then divide that number by 1000 to get it into kg, then multiply by 2.2. \[[g] \cdot \left [1 kg \over 1000 g \right] \cdot \left [ 2.2 lbs \over 1 kg \right]\]

OpenStudy (anonymous):

would the answer be 59.6 pounds?

OpenStudy (anonymous):

If you have the volume in Liters, the dimensional analysis will be\[\rho \left [g \over cm^3 \right] \cdot V \left [ L \right] \cdot \left [1000 cm^3 \over 1L \right] \cdot \left [1 kg \over 1000 g \right] \cdot \left [ 2.2 lbs \over 1 kg \right ] = mass [ lbs]\]

OpenStudy (anonymous):

I got 59.4 lbs

OpenStudy (anonymous):

oh right i checked my math and I got 59.4, thank you so much for your help!

OpenStudy (anonymous):

Welcome.